Find killing equations for spherical polar coordinates


1
Expert's answer
2021-12-29T16:30:21-0500

Killing vector field is a vector field on a Riemannian manifold that preserves the metric.

Killing equation:

"\nabla_{\mu}X_{v}+\nabla_{v}X_{\mu}=0"

Killing fields on a 2-sphere:

The conventional metric on the sphere is

"g(\Omega)=d\theta^2+sin^2\theta d\phi^2"

Rotation about the pole should be an isometry. The infinitesimal generator of a rotation can then be identified as a generator of the Killing field. This can be

"U=\partial_ \phi"

The surface of the sphere is two-dimensional, and there is another generator of isometries; it can be taken as

"V=\partial_ \theta"


conventional 3-space coordinate system is given by

"x=sin\theta,y=sin\theta sin \phi,z=cos\theta"

first generator, rotation about the  z-axis:

"R=x\partial _y-y\partial_x=sin^2\theta \partial \phi"

second generator, rotation about the  x-axis:

"S=z\partial _y-y\partial_z"

third generator, rotation about the  x-axis:

"T=z\partial _x-x\partial_z"


Normalizing this, and expressing these in spherical coordinates:


"R'=\frac{R}{sin^2\theta}=\partial_ \phi"


"S'=\frac{S}{sin^2\theta}=sin\phi \partial_ \theta +cot\theta cos\phi \partial _\phi"


"T'=\frac{T}{sin^2\theta}=cos\phi \partial_ \theta +cot\theta sin\phi \partial _\phi"